OpenStudy (anonymous):

Juan deposited $200 in a savings account earning 3% interest over 2 years. What was the total amount of interest earned at the end of 2 years?

undeadknight26 (undeadknight26):

hmmm f(x) = 200(1.3)^x? or maybe.. f(x) = 200(1.3)x?